OpenStudy (marco26):

Need help. Step by step. Find the differential equation of all circles tangent to the x-axis.

OpenStudy (marco26):

@anusha.p . After knowing (x-a)^2 + (y-b)^2 = b^2, what's next? Differentiate? Thanks

OpenStudy (marco26):

Differentiating wrt x will give: 2(x-a)+2(y-b)y' = 0 (x-a)+(y-b)y"=0 Then, how to differentiate wrt y?

OpenStudy (marco26):

The answer is supposed to be like this: [1+(y')^2]^3= [yy''+(y')^2]^2. Will it look something like that?

OpenStudy (anonymous):

i did it but i get [1+(y')^2]^3= [1+yy''+(y')^2]^2 i dont know why let me check it again

OpenStudy (anonymous):

i think my answer is correct let me show my work give me a sec

OpenStudy (marco26):

ok

OpenStudy (anonymous):

here

OpenStudy (marco26):

@mukushla thanks! i get it now,,can i ask one more problem?

OpenStudy (anonymous):

yes. but if it is math plz post it in math category. tnx

OpenStudy (marco26):

how about find the differential equation of all straight lines unit distance from the origin.

OpenStudy (marco26):

@mukushla